OpenStudy (anonymous):

Demonstrate how a recursive process will allow you to find the number of coins and points on all levels up to level 5. @texaschic101 @Jonathan1997 Last question sorry

OpenStudy (anonymous):

I know its a rule that can be repeated

OpenStudy (anonymous):

I think I got it A recursive process will allow me to use the rule of plus 2 and *3 repeatedly until i get the amount coins and points to get to level 5 and the amount of coins.

OpenStudy (anonymous):

if you write something i will give you a medal for coming when i asked

OpenStudy (anonymous):

an = 15(3)^x is this a sequence formula for the amount of points on level 15 with a common ratio of 3?

OpenStudy (anonymous):

Sorry this an = 15(2*3)^x

OpenStudy (anonymous):

common ratio of 2 and common ratio of 3

OpenStudy (anonymous):

Texas? are you even actually typing or are you just typing a bunch of stuff.

OpenStudy (texaschic101):

a recursive formula is a formula that requires the computation of all previous terms in order to find the value of an. example : a1 = 3 an = 2a (n-1) + 5 a2 = 2(a1) + 5 = 2(3) + 5 = 11 a3 = 2(a2) + 5 = 2(11) + 5 = 27 a4 = 2(a3) + 5 = 2(27) + 5 = 59 and so on... as far as your 15(3)^x is a sequence formula.....I think I am not the best at this

OpenStudy (anonymous):

okay so my sequence formula is good but my recursive formula needs work.

OpenStudy (texaschic101):

you would probably need to ask somebody that knows more about this. I am so sorry.

OpenStudy (anonymous):

it was helpful though

OpenStudy (texaschic101):

typo * should

OpenStudy (anonymous):

could you help me make my recursive formula if i have two common ratios one +2 and the other *3 and 15 levels

OpenStudy (texaschic101):

it would be better if you got somebody who is better at this.....I am so sorry

OpenStudy (anonymous):

Jonothan?

OpenStudy (texaschic101):

possibly ?

OpenStudy (anonymous):

@dan815

OpenStudy (texaschic101):

good luck :)

OpenStudy (anonymous):

IM calling the calvary!

OpenStudy (texaschic101):

@satellite73

OpenStudy (texaschic101):

@Directrix ...can you help

OpenStudy (dan815):

can you tell me what your question is exactly

OpenStudy (dan815):

its all in parts up there

OpenStudy (anonymous):

3. The development team has asked you to jump ahead of them in the project. Create the sequence formulas, an, for the coins and the points based on the level in the game. Then describe how the formula can be used to find the coins and values on level 15. Use complete sentences.

OpenStudy (anonymous):

i got I use 15 because of the level and 2*3 for the common ratios which are 2 and 3. an = 15(2*3)^x which i believe to be wrong.

OpenStudy (dan815):

okay and how do these coins build up a levl

OpenStudy (anonymous):

Level Coins 1 2 2 4 3 6 Level Points 1 3 2 9 3 27

OpenStudy (anonymous):

common ratio for coins +2 common ratio for points *3

OpenStudy (dan815):

okay so level and coins have +2 relation and levl and points have 3^level relation

OpenStudy (dan815):

u want sequence formula or recursive?

OpenStudy (anonymous):

ok so an = 2(3)^15

OpenStudy (anonymous):

sequence

OpenStudy (dan815):

okay sequence is 3^15 for 15 levels

OpenStudy (dan815):

3^15= points at 15 2*15= coins at 15

OpenStudy (anonymous):

ok one more part which I believe to have right 4. If the game only has 20 levels, explain how to find the value of the series for the coins and the points. Use complete sentences and arrive at final values. You would use You could just keep multiplying the points by 3 and adding the points by 2 until you get 20. Which would result in in 40 coins equaling 3486784401points. (kinda outrageous for only 40 coins but i believe the math is correct.)

OpenStudy (dan815):

i guess so i dont know what 3^20 is, its definately a huge number

OpenStudy (anonymous):

also what would that make the sequence equation 15(2)3^15 I know this is incorrect but I don't know where to get with it.

OpenStudy (dan815):